OpenStudy (anonymous):

determine whether y varies directly with x. if so, find the constant of variation k. -4y=8x+16

OpenStudy (anonymous):

it does and just solve for y and you will get it in y=mx+b m would be your k i think.

OpenStudy (ganpat):

-4y=8x+16 y = -2x - 8 so y = - (2x + 8).. since its negative, its inversely proportional...

OpenStudy (anonymous):

so would my k be 2? or 8...

OpenStudy (anonymous):

2 your m

OpenStudy (ganpat):

i am not sure.. i guess, it should be 2...

OpenStudy (anonymous):

thanks for the help!!

OpenStudy (ganpat):

w/c...